OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 2x + 9y = 4 3x + 7y = –7

OpenStudy (anonymous):

2x + 9y = 4 ...............1 mul. by 3 3x + 7y = –7................2 mul. by -2 6x +27y=12...............1 -6x -14y=14...............2 ---------------- 0 +13 y = 26 y=2 sub y in equation 1 2x+(9×2)=4 ----> 2x=4-18 x=-7